"Three organisms were innoculated on a culture media. Results showed Escherichia coli as red colonies, no growth of Staphylococcus epidermidis, and colorless colonies of Salmonella enterica. The medium is:"

a

selective and differential

b

differential

c

selective

d

minimal

e

heterotrophic

"A bacterium was incubated with aeration in a nutrient medium containing two carbon sources. After logarithmic growth, a lag phase was observed for 25 minutes. Then logarithmic growth at a slower rate was observed. Which substate was preferred by the bacteria?"

a

The first substrate used up is preferred.

b

The second substrate used up is preferred.

c

Neither substate is preferred.

d

Both substrates are preferred equally.

e

not enough information is given to determine the correct answer

"A culture of 10 bacteria of Escherichia coli has a generation time of 30 minutes at 35 degrees Celsius. After 5 hours at this temperature, how many bacteria would you expect to see?"

a

"10,240"

b

50

c

"1,000"

d

320

e

"10,000,000,000"

Pasteurization of 63 degress Celsius for 30 minutes or 72 degrees Celsius for 15 seconds best illustrate

a

equivalent treatments

b

methods of sterlization

c

methods of filtration

d

how thermoduric bacteria survive

e

decimal reduction time

Why is Mycobacterium often resistant to disinfectants?

a

acid-resistant cell wall

b

endospore formation

c

metabolism

d

anaerobic behavior

e

genetic recombination

Microwave radiation was examined in wet and dry conditions. What is the most likely method of lethal action of microwave radiation?

a

heat

b

ionizing radiation

c

pressure

d

moisture

e

all of the choices are correct

Which causes protein denaturation and is effective against enveloped viruses?

a

alcohols

b

halogens

c

organic acids

d

biguanides

e

silvera
Answer #1

Ans-1- selective because kills E Coli and differential because help to differentiate between Staphylococcus epidermidis and Salmonella enterica

So the option A is correct

Ans-2- Since the rate of growth in second phase is slow the first carbon source is preferred by bacteria.

So the option A is correct

Ans-3- N0 =10 G.T. = 30 min Total time = 5 hrs = 300 min.

So the number of generation (n) = 300 / 30 =10

from the formula

N = N0 * 2n = 10 * 210 = 10240 cells

So the option A is correct

Ans-4- Both the treatment are equivalent

So the option A is correct

Ans-5- mycobacterium contains an acid-resistant cell wall, that contains a waxy layer and prevents disinfectant to enter into the cell

So the option A is correct

Ans-6- Microwave heat the bacteria that cause loss of water and due dehydration bacteria dies.

So the option A is correct

Ans-7- Alcohol is most commonly used disinfactant that kills pathogen by denauring their proteins

So the option A is correct